OpenStudy (anonymous):

What is the exact value of the expression Sqrt(18) + sqrt( 48) - sqrt(50)? Simplify completely.

OpenStudy (earthcitizen):

\[4 \sqrt(3)-2 \sqrt(2)\]

OpenStudy (anonymous):

√18 + √48 - √50 = 3√2 + 4√3 - 5√2 = 4√3 - 2√2

OpenStudy (anonymous):

I appreciate it!! I'm a little confused....these were my options sqrt(16) -14 sqrt(2) + 16 sqrt(3) -2 sqrt(2) + 4 sqrt(3) - sqrt(8) + sqrt (48)

OpenStudy (anonymous):

The third one

OpenStudy (anonymous):

You just square what's outside the radical and put it back into the radical.

OpenStudy (anonymous):

ahhh! Thanks!:) I am not good at these types of things! Thanks for helping me.

OpenStudy (anonymous):

4√3 - 2√2 = √16*3 - √4*2 = √48 - √8 = -√8 + √48